OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. -13 4 13 -12

OpenStudy (anonymous):

(-3-6)/(10-x) =-3/2 9/(10-x)=3/2 6=10-x x=4

OpenStudy (anonymous):

simply use the slope formula(y2-y1)/(x2-x1)

OpenStudy (anonymous):

slope of a line which passes through given two points =(Y1-Y2)/(X1-X2) -3/2=(6-(-3))/(x-10)=>x=4

OpenStudy (anonymous):

you use either (Y1-Y2)/(X1_X2) or (Y2-Y1)/(X2-X1)